# Searching in Euclidean Spaces with Predictions

### v1

@@ 4.1 Extensions
First, the same approach works for spaces of bounded doubling dimension as long as there is a concept of path to connect points whose length is the same as the distance or we change the setting so that the cost of moving from one point to another is the distance between the points. For this, one has to use $\varepsilon$-nets in spaces of bounded doubling dimension [18].
When the target is known to have integral coordinates, we can modify the search so that it has a finite number of segments. Indeed, as soon as we reach a point $p_{i}$ such that $\lambda(p_{i})<1/2$, we know that the target is at distance smaller than $1/2$ from $p_{i}$ and there is a unique point with integral coordinates in $B(p_{i},\lambda(p_{i}))$. We can then just move that point. A similar approach works when the target is known to have coordinates with bounded resolution by scaling the setting. In the case where we can detect the target when we are within a given distance $\delta$, we can finish the search when we reach a point $p_{i}$ with $\lambda(p_{i})\leq\delta$, thus, also bounding the number of steps (which will depend on the initial estimate $\lambda(p_{0})$).
@@ 5 Lower bound
In this section we provide a lower bound for the search problem with $c$-predictions. Our lower bounds are meaningful when we assume that $c$ is large enough. The idea is to construct many $c$-predictions for several different targets that are indistinguishable, unless we are quite close to the target. For each such prediction $\lambda$, there is a small ball around the target where the value of $\lambda$ is different from the other predictions, but all the other predictions have the same value on that small ball. Then, the searcher has to visit all the small balls around the targets, because in the worst case the target is going to be in the last small ball visited.
First, using the techniques by Elbassioni, Fishkin and Sitters [12] developed for the Euclidean Traveling Salesperson with Neighbourhoods, one can show the following bound. See Dumitrescu and Tóth [11] for an improvement over [12] where the same idea is reused.
@@
Let $\mathbb{B}$ be a set of $n$ congruent balls of radius $\delta$ in $\mathbb{R}^{d}$ that are pairwise interior disjoint. Each path in $\mathbb{R}^{d}$ that contains at least one point from each ball of $\mathbb{B}$ has length at least $\big{(}n/2^{d}-1\big{)}\cdot\delta\cdot\sqrt{\frac{\pi}{d}}$.
@@
Note that for any ball $B_{i}\in\mathbb{B}$ and any point $p\in B_{i}$, the ball $B(p,\delta)$ intersects $B_{i}$ in a large volume. More precisely, because $B_{i}\cap B(p,\delta)$ contains a ball of radius $\delta/2$, we have
$\forall p\in B_{i}:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{vol}\big{(}B_{i}\cap B(p,\delta)\big{)}\geq V_{d}(\delta/2)=(\delta/2)^{d}\cdot V_{d}(1).$
Consider any path $\pi$ that touches each ball of $\mathbb{B}$ and let $L$ be its length. Consider the volume of the Minkowski sum $\pi\oplus B(\bm{o},\delta)$ of the (points on the) path $\pi$ and the ball $B(\bm{o},\delta)$, that is the set of points at distance at most $\delta$ from $\pi$. Since the path touches each ball $B_{i}$ of $\mathbb{B}$ at some point $p_{i}\in B_{i}$, and the balls in $\mathbb{B}$ are interior pairwise disjoint, we have
$\operatorname{vol}\big{(}\pi\oplus B(\bm{o},\delta)\big{)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{B_{i}\in\mathbb{B}}\operatorname{vol}\big{(}B_{i}\cap B(p_{i},\delta)\big{)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ n\cdot(\delta/2)^{d}\cdot V_{d}(1).$
@@
$L\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{n\cdot(\delta/2)^{d}\cdot V_{d}(1)-\delta^{d}\cdot V_{d}(1)}{\delta^{d-1}\cdot V_{d-1}(1)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{(}n/2^{d}-1\big{)}\cdot\delta\cdot\frac{V_{d}(1)}{V_{d-1}(1)}.$
@@
(For $d=1$ we can just check that $\Gamma(1)=1>\frac{\sqrt{\pi}}{2}=\Gamma(3/2)$). Thus, $L\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \big{(}n/2^{d}-1\big{)}\cdot\delta\cdot\sqrt{\frac{\pi}{d}}$.
Next, fix a value $c>2$. For each $\bm{t}\in B(\bm{o},1/2-1/c)$, let $\lambda_{\bm{t}}$ be the function defined as follows; see Figures 5 and 6 for intuition.
$\lambda_{\bm{t}}(p)=\begin{cases}c\cdot|p\bm{t}|,&\text{ if $x\in B(\bm{t},1/c)$},\\ 1,&\text{ if $x\in B(\bm{o},1/2)\setminus B(\bm{t},1/c)$},\\ 2\cdot|p\bm{o}|,&\text{ if $x\notin B(\bm{o},1/2)$}.\end{cases}$ (2)
@@
Assume that $c>2$ and $|\bm{o}\bm{t}|\leq\tfrac{1}{2}-\tfrac{1}{c}$. Then the function $\lambda_{\bm{t}}$ is a $c$-prediction for the target $\bm{t}$. Moreover, for any two distinct $\bm{t},\bm{t}^{\prime}$ satisfying the hypothesis, the functions $\lambda_{\bm{t}}$ and $\lambda_{\bm{t}^{\prime}}$ agree on all points outside $B(\bm{t},1/c)\cup B(\bm{t}^{\prime},1/c)$.
@@
Consider any fixed $\bm{t}\in S(\bm{o},1/c,1/2)$. We have to show that $\lambda_{\bm{t}}$ satisfies
@@
Assume that $c\geq 4$. There is family of $c$-predictions for targets in $B(\bm{o},1/4)$ such that any search path in $\mathbb{R}^{d}$ that uses $c$-predictions to find the target has length at least $(c^{d-1}/16^{d})\cdot\min\{\sqrt{\pi/d},1\}$.
@@
Set $\varepsilon=2/c$. Because of Lemma 2.2, the ball $B(\bm{o},1/4)$ has a $\varepsilon$-separated set $N$ with
$|N|\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \left(\frac{1/4}{2/c}\right)^{d}\leavevmode\nobreak\ =\leavevmode\nobreak\ (c/8)^{d}$
points. Note that the balls $B(q,1/c)$, where $q\in N$, are pairwise interior disjoint.
@@
For each point $\bm{t}\in N$, consider the function $\lambda_{\bm{t}}$ defined in equation (2). Note that $\frac{1}{4}\leq\frac{1}{2}-\frac{1}{c}$ because $c\geq 4$, and therefore $B(\bm{o},1/4)\subseteq B(\bm{o},1/2-1/c)$, as required for the definition. Lemma 5.3 shows that the function $\lambda_{\bm{t}}$ is a $c$-prediction for each $\bm{t}\in N$. Consider the set of prediction functions $\Lambda=\{\lambda_{\bm{t}}\mid\bm{t}\in N\}$. Let $\mathbb{B}=\{B(\bm{t},1/c)\mid\bm{t}\in N\}$ and note that these balls are pairwise interior disjoint. All the functions $\lambda\in\Lambda$ take the same value on each point $p$ outside $\bigcup\mathbb{B}$. Moreover, to distinguish between two functions $\lambda_{\bm{t}},\lambda_{\bm{t}^{\prime}}\in\Lambda$ we have to evaluate them at some point of $B(\bm{t},1/c)\cup B(\bm{t}^{\prime},1/c)$.
@@
Consider any search strategy when the $c$-prediction function is selected from $\Lambda$ by an adversary. It may use that the $c$-prediction is from $\Lambda$. In particular, we know beforehand that the target $\bm{t}$ is a point of $N$, selected by the adversary. We claim that the search path has to visit the interior of all the balls of $\mathbb{B}$. Indeed, while there are two distinct points $\bm{t},\bm{t}^{\prime}\in N$ such that the balls $B(\bm{t},1/c)$ and $B(\bm{t}^{\prime},1/c)$ are not visited by the search path, we have $\lambda_{\bm{t}}(p)=\lambda_{\bm{t}^{\prime}}(p)$ for all points $p$ along the path, and therefore the information collected cannot discern whether the target is $\bm{t}$ or $\bm{t}^{\prime}$. Thus, in the worst case, the search path has to visit all the balls of $\mathbb{B}$ but one, to identify which point of $N$ is the target, and then still may have to move to that last ball. That is, we may assume that in the worst case the adversary chooses $\lambda_{\bm{t}}\in\Lambda$ such that $B(\bm{t},1/c)$ is the last ball of $\mathbb{B}$ visited by the search strategy. After the searcher deduces at which point of $N$ the target is, they still have to move there.
@@
We conclude that the search path has to visit all $|N|\geq(c/8)^{d}$ balls of $\mathbb{B}$, and for the last ball it still has to travel $1/c$ to the center. Since those balls are pairwise interior disjoint, Lemma 5.1 implies that the search path has length at least
$\frac{1}{c}+\left(\frac{|N|}{2^{d}}-1\right)\cdot\left(\frac{1}{c}\right)\cdot\sqrt{\frac{\pi}{d}}\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \frac{1}{c}+\left(\frac{|N|}{2^{d}}-1\right)\cdot\left(\frac{1}{c}\right)\min\{\sqrt{\pi/d},1\}$
$\geq\leavevmode\nobreak\ \frac{|N|}{2^{d}}\cdot\left(\frac{1}{c}\right)\cdot\min\{\sqrt{\pi/d},1\}$
$\geq\leavevmode\nobreak\ \frac{(c/8)^{d}}{2^{d}}\cdot\left(\frac{1}{c}\right)\cdot\min\{\sqrt{\pi/d},1\}$
$=\leavevmode\nobreak\ \frac{c^{d-1}}{16^{d}}\cdot\min\{\sqrt{\pi/d},1\}.$
@@
For the competitive ratio we have to compare the length of the search path to the distance to the target, which in the construction is at most $1/4$. We then obtain the following bound, which is interesting when $d\geq 2$ and $c$ is large enough.
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ with prediction factor $c\geq 4$. Any search strategy to reach the target has competitive ratio at least $(1/4)\cdot(c/16)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$.
@@
Consider the construction of Theorem 5.5. Since the search path has length at least $(c^{d-1}/16^{d})\cdot\min\{\sqrt{\pi/d},1\}$ and the target is $\bm{t}\in B(\bm{o},1/4)$, the competitive ratio is at least
$\frac{(c^{d-1}/16^{d})\cdot\min\{\sqrt{\pi/d},1\}}{|\bm{o}\bm{t}|}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{(c^{d-1}/16^{d})\cdot\min\{\sqrt{\pi/d},1\}}{1/4}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{4}\cdot\frac{c^{d-1}}{16^{d-1}}\cdot\min\{\sqrt{\pi/d},1\}.$
The lower bound still holds in the case where the target has integral coordinates or is detected when we are within a given distance, with a slightly worse constant in the basis of the exponential function. For this, we can just scale the input such that the balls in the construction are large enough or the epsilon-net can be assumed to have points with integer coordinates. Then, we still have to visit all balls in order to detect which one has the target.
A similar lower bound holds for randomized search strategies. In our construction, the searcher has to visit a certain number of disjoint regions (balls) and they can deduce something only when visiting the ball that contains the target. In the worst-case (or adversary) model, the target is going to be in the last ball that is visited. In a randomized setting, all balls would be equally likely to contain the target, and thus the searcher has to visit in expectation half of the balls because visiting the wrong ball does not provide any information. Thus, we have a TSP where you have to visit (in expectation) half of the pairwise disjoint balls, and a similar lower bound applies because the bound is linear in the number of balls.
6 Conclusions and open problems
We have introduced and studied the problem of searching for a target in $\mathbb{R}^{d}$ under a predictions model in which the searcher is given, at each position they visit, an approximate distance to the target. We presented strategies with competitive ratio $(6c)^{d+1}/3$ when the prediction factor $c$ is known, and $(12c)^{d+1}$ when the prediction factor $c$ is unknown. We also showed a lower bound of $(1/4)\cdot(c/16)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$ for the competitive ratio of any strategy under this model, even when the prediction factor $c$ is known.

### v2

@@ 4.1 Extensions
First, the same approach works for spaces of bounded doubling dimension as long as there is a concept of path to connect points such that the length of the path is the same as the distance between the points. We could also change the setting so that the cost of moving from one point to another is the distance between the points. For this, one has to use $\varepsilon$-nets in spaces of bounded doubling dimension [19].
When the target is known to have integral coordinates, we can modify the search so that it has a finite number of segments. Indeed, as soon as we reach a point $p_{i}$ such that $\lambda(p_{i})<1/2$, we know that the target is at distance smaller than $1/2$ from $p_{i}$ and there is a unique point with integral coordinates in $B(p_{i},\lambda(p_{i}))$. We can then just move to that point. A similar approach works when the target is known to have coordinates with bounded resolution by scaling the setting. In the case where we can detect the target when we are within a given distance $\delta$, we can finish the search when we reach a point $p_{i}$ with $\lambda(p_{i})\leq\delta$, thus, also bounding the number of steps (which will depend on the initial estimate $\lambda(p_{0})$).
@@ 5 Lower bound
In this section we provide a lower bound for the search problem with $c$-predictions. Our lower bounds are meaningful when we assume that $c$ is large enough. The idea is to construct $c$-predictions for infinitely many different targets with the property that any two of them are indistinguishable, unless we are quite close to the target. For each such prediction $\lambda$, there is a small ball around the target where the value of $\lambda$ is different from the other predictions, but most of the other predictions have the same value on that small ball. Then, the searcher has to visit all the small balls around the targets, because in the worst case the target is going to be in the last small ball that is visited.
First, inspired by the technique by Elbassioni, Fishkin and Sitters [13] developed for the Euclidean Traveling Salesperson with Neighbourhoods, one can show the following bound. See Dumitrescu and Tóth [12] for an improvement over [13] where the same idea is reused.
@@
For any given radius $\delta\leq 1/2$ and dimension $d\geq 1$, consider the infinite family of balls
$\mathbb{B}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{\{}B(p,\delta)\subset\mathbb{R}^{d}\mid p\in B(\bm{o},1/2-\delta)\big{\}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{\{}B(p,\delta)\subset\mathbb{R}^{d}\mid B(p,\delta)\subseteq B(\bm{o},1/2)\big{\}}.$
Each path in $\mathbb{R}^{d}$ that contains at least one point from each ball of $\mathbb{B}$ has length at least $\left(\big{(}\frac{1}{2\delta}-1\big{)}^{d}-1\right)\cdot\delta\cdot\sqrt{\frac{\pi}{d}}$.
@@
Consider any path $\pi$ that touches each ball of $\mathbb{B}$ and let $L$ be its length. Consider the volume of the Minkowski sum $\pi\oplus B(\bm{o},\delta)$ of the (points on the) path $\pi$ and the ball $B(\bm{o},\delta)$, that is the set of points at distance at most $\delta$ from $\pi$. Since the path touches each ball of $\mathbb{B}$, the set $\pi\oplus B(\bm{o},\delta)$ contains the center of each ball of $\mathbb{B}$. This means that $\pi\oplus B(\bm{o},\delta)$ contains $B(\bm{o},1/2-\delta)$, and therefore we have
$\operatorname{vol}\big{(}\pi\oplus B(\bm{o},\delta)\big{)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \operatorname{vol}\big{(}B(\bm{o},1/2-\delta)\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \left(\frac{1}{2}-\delta\right)^{d}\cdot V_{d}(1).$
@@
$L\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\left(\frac{1}{2}-\delta\right)^{d}\cdot V_{d}(1)-\delta^{d}\cdot V_{d}(1)}{\delta^{d-1}\cdot V_{d-1}(1)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \left(\Big{(}\frac{1}{2\delta}-1\Big{)}^{d}-1\right)\cdot\delta\cdot\frac{V_{d}(1)}{V_{d-1}(1)}.$
@@
(For $d=1$ we can just check that $\Gamma(1)=1>\frac{\sqrt{\pi}}{2}=\Gamma(3/2)$.) The result follows.
Next, fix a value $c\geq 2$. For each $\bm{t}\in B(\bm{o},1/2-1/c)$, let $\lambda_{\bm{t}}$ be the function defined as follows; see Figures 5 and 6 for intuition.
$\lambda_{\bm{t}}(p)=\begin{cases}c\cdot|p\bm{t}|,&\text{ if $p\in B(\bm{t},1/c)$},\\ 1,&\text{ if $p\in B(\bm{o},1/2)\setminus B(\bm{t},1/c)$},\\ 2\cdot|p\bm{o}|,&\text{ if $p\notin B(\bm{o},1/2)$}.\end{cases}$ (2)
@@
Assume that $c\geq 2$ and $|\bm{o}\bm{t}|\leq\tfrac{1}{2}-\tfrac{1}{c}$. Then the function $\lambda_{\bm{t}}$ is a $c$-prediction for the target $\bm{t}$. Moreover, for any two distinct $\bm{t},\bm{t}^{\prime}$ satisfying the hypothesis, the functions $\lambda_{\bm{t}}$ and $\lambda_{\bm{t}^{\prime}}$ agree on all points outside $B(\bm{t},1/c)\cup B(\bm{t}^{\prime},1/c)$.
@@
Consider any fixed $\bm{t}\in B(\bm{o},1/2-1/c)$. We have to show that $\lambda_{\bm{t}}$ satisfies
@@
Assume that $c>2$. There is family of $c$-predictions for targets in $B(\bm{o},1/2-1/c)$ such that any search path in $\mathbb{R}^{d}$ that uses $c$-predictions to find the target has length at least $\left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}$.
@@
For each $\bm{t}\in B(\bm{o},1/2-1/c)$, consider the function $\lambda_{\bm{t}}$ defined in equation (2). Lemma 5.3 shows that the function $\lambda_{\bm{t}}$ is a $c$-prediction for each $\bm{t}\in N$. Consider the set of prediction functions $\Lambda=\{\lambda_{\bm{t}}\mid\bm{t}\in B(\bm{o},1/2-1/c)\}$. Let $\mathbb{B}=\{B(\bm{t},1/c)\mid\bm{t}\in B(\bm{o},1/2-1/c)\}$. Note that to distinguish between two functions $\lambda_{\bm{t}},\lambda_{\bm{t}^{\prime}}\in\Lambda$ we have to evaluate them at some point of $B(\bm{t},1/c)\cup B(\bm{t}^{\prime},1/c)$.
@@
Consider any search strategy when the $c$-prediction function is selected from $\Lambda$ by an adversary. It may use that the $c$-prediction is from $\Lambda$. In particular, we know beforehand that the target $\bm{t}$ is a point of $B(\bm{o},1/2-1/c)$, selected by the adversary. We claim that the search path has to visit all the balls of $\mathbb{B}$. Indeed, while there are two distinct points $\bm{t},\bm{t}^{\prime}\in B(\bm{o},1/2-1/c)$ such that the balls $B(\bm{t},1/c)$ and $B(\bm{t}^{\prime},1/c)$ are not visited by the search path, we have $\lambda_{\bm{t}}(p)=\lambda_{\bm{t}^{\prime}}(p)$ for all points $p$ along the path, and therefore the information collected cannot discern whether the target is $\bm{t}$ or $\bm{t}^{\prime}$. Thus, in the worst case, the search path has to visit all the balls of $\mathbb{B}$ but one, to identify which point of $B(\bm{o},1/2-1/c)$ is the target, and then still may have to move to that last ball. That is, we may assume that in the worst case the adversary chooses $\lambda_{\bm{t}}\in\Lambda$ such that $B(\bm{t},1/c)$ is the last ball of $\mathbb{B}$ visited by the search strategy. After the searcher deduces at which point of $B(\bm{o},1/2-1/c)$ the target is, they still have to move to the target.
@@
We conclude that the search path has to visit all balls of $\mathbb{B}$, and for the last ball it still has to travel $1/c$ to the center. Lemma 5.1 (with $\delta=1/c$) implies that the search path has length at least
$\frac{1}{c}+\left(\Big{(}\frac{c}{2}-1\Big{)}^{d}-1\right)\cdot\frac{1}{c}\cdot\sqrt{\frac{\pi}{d}}\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \frac{1}{c}+\left(\Big{(}\frac{c-2}{2}\Big{)}^{d}-1\right)\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}$
$\geq\leavevmode\nobreak\ \left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}.$
@@
For the competitive ratio we have to compare the length of the search path to the distance to the target, which in the construction is at most $1/2-1/c$. We then obtain the following bound, which is interesting when $d\geq 2$ and $c>2$.
@@
Consider the search with predictions problem in $\mathbb{R}^{d}$ with prediction factor $c>2$. Any search strategy to reach the target has competitive ratio at least $\left(\frac{c-2}{2}\right)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$.
@@
Consider the construction of Theorem 5.5. Since the search path has length at least $\left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}$ and the target is $\bm{t}\in B(\bm{o},1/2-1/c)$, the competitive ratio is at least
$\frac{\left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}}{|\bm{o}\bm{t}|}$ $\geq\leavevmode\nobreak\ \frac{\left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}}{1/2-1/c}$
$=\leavevmode\nobreak\ \frac{\left(\frac{c-2}{2}\right)^{d}\cdot\frac{1}{c}\cdot\min\{\sqrt{\pi/d},1\}}{\frac{c-2}{2c}}$
$=\leavevmode\nobreak\ \left(\frac{c-2}{2}\right)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}.$
To get a bound that is easier to grasp, let us assume that $c\geq 4$. In this case $(c-2)/2\geq c/4$ and we get the following bound.
Corollary 5.9.
Consider the search with predictions problem in $\mathbb{R}^{d}$ with prediction factor $c\geq 4$. Any search strategy to reach the target has competitive ratio at least $(c/4)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$.
The lower bounds still hold in the case where the target has integral coordinates or is detected when we are within a given distance, with a slightly worse constant in the basis of the exponential function. For this, we can just scale the input such that the balls in the construction are large enough to contain points with integer coordinates. Then, we still have to visit all balls in order to detect which one has the target.
A similar lower bound holds for randomized search strategies. In our construction, the searcher has to visit some balls in the family $\mathbb{B}$ of balls of Lemma 5.1 (with $\delta=1/c$), and they can deduce something only when visiting the ball whose center is the target. In the worst-case (or adversary) model, the target is going to be the center of the last ball of $\mathbb{B}$ that is visited. In a randomized setting, we select the target uniformly at random among all the possible ones. This means that $\bm{t}$ is selected uniformly at random among the points of $B(\bm{o},1/2-1/c)$. When the search path has touched half of the balls of $\mathbb{B}$ (in a measure theoretic sense), with probability $1/2$ the ball $B(\bm{t},1/c)$ has been touched. A similar argument as the one used in the proof of Lemma 5.1 implies that any search path $\pi$ that touches half of the balls of $\mathbb{B}$ must satisfy the volume inequality
$\operatorname{vol}\big{(}\pi\oplus B(\bm{o},\delta)\big{)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}\cdot\operatorname{vol}\big{(}B(\bm{o},1/2-\delta)\big{)}.$
With this we obtain a similar lower bound for randomized search strategies, where the competitive ratio is about $1/4$ worse.
6 Conclusions and research directions
We have introduced and studied the problem of searching for a target in $\mathbb{R}^{d}$ under a predictions model in which the searcher is given, at each position they visit, an approximate distance to the target. We presented strategies with competitive ratio $4\cdot 5^{d}\cdot c^{d+1}$ when the prediction factor $c$ is known, and $6\cdot 10^{d}\cdot c^{d+1}$ when the prediction factor $c$ is unknown. We also showed a lower bound of $(c/4)^{d-1}\cdot\min\{\sqrt{\pi/d},1\}$ for the competitive ratio of any strategy under this model, assuming that $c\geq 4$, even when the prediction factor $c$ is known.